OpenStudy (anonymous):

Show the expanded form and simplification of x to the 6th power over x to the second power. Explain in your own words how you can simplify x to the 6th power over x to the second power without having to write the expanded form. Part 2: Create your own fraction with like bases, coefficients, and show its simplification.

OpenStudy (phi):

I assume you know that in \( x^6\) the little 6 in the upper right means multiply x times itself 6 times? It is short-hand, and helpful for x^20

OpenStudy (jdoe0001):

$$ \cfrac{x^6}{x^2} \implies \cfrac{x^4}{1} \times\cfrac{x^2}{x^2}\\ \color{blue}{\cfrac{x^2}{x^2} = 1}\\ \cfrac{x^4}{1} \times\cfrac{x^2}{x^2} \implies \cfrac{x^4}{1} \times 1 \implies x^4 $$

OpenStudy (phi):

expanded form probably means write out x times itself 6 times. can you count to six ?

OpenStudy (anonymous):

lol yes

OpenStudy (jdoe0001):

$$\large{ \cfrac{x^6}{x^2} \implies \cfrac{x}{1} \times\cfrac{x}{1}\times\cfrac{x}{1}\times\cfrac{x}{1}\times\cfrac{x}{x}\times\cfrac{x}{x} } $$

OpenStudy (phi):

to simplify you need to know anything divided by itself is 1 so x/x is 1 and 1 times anything is anything (so you can ignore the 1 if you are multiplying)

OpenStudy (anonymous):

okay so its x^4

OpenStudy (phi):

it is x*x*x*x which can also be written as \( x^4\)

OpenStudy (anonymous):

okay thanks!